# Towards Imperceptible Backdoor Attack in Self-supervised Learning

### v1

@@ Abstract
Self-supervised learning models are vulnerable to backdoor attacks. Existing backdoor attacks that are effective in self-supervised learning often involve noticeable triggers, like colored patches, which are vulnerable to human inspection. In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models. We first find that existing imperceptible triggers designed for supervised learning are not as effective in compromising self-supervised models. We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in self-supervised learning. Building on this insight, we design an attack using optimized triggers that are disentangled to the augmented transformation in the self-supervised learning, while also remaining imperceptible to human vision. Experiments on five datasets and seven SSL algorithms demonstrate our attack is highly effective and stealthy. It also has strong resistance to existing backdoor defenses. Our code can be found at https://github.com/Zhang-Henry/IMPERATIVE.
@@ 1 Introduction
In recent years, Self-Supervised Learning (SSL) has emerged as a powerful paradigm in machine learning, particularly within the computer vision domain. The key advantage of SSL lies in its ability to learn abundant representations from the vast amounts of unlabeled data, bypassing the labor-intensive and costly process of manual labeling. The objective of SSL in computer vision is to develop an image encoder, designed to ensure that similar images yield close embeddings in the encoder. This is achieved by constructing similar images by adding different augmentations on the same image. Utilizing this pre-trained encoder, users can conduct different downstream tasks by training downstream classifiers based on it. The downstream classifiers are typically compact networks with a small number of parameters.
@@
Although SSL has been extensively used in the development of foundational models [5, 22, 4], it is at risk of backdoor attacks [29, 61, 37, 75], where the attacker embeds hidden malicious behavior within the encoder. The backdoor can be inherited to the downstream task. That is, the downstream classifier will predict a specific target label if the input contains a pre-defined backdoor trigger. Existing backdoor attacks on SSL such as BadEncoder [30] achieve high attack success rates. However, a common drawback of these effective attacks is that their trigger patterns are obvious, making them susceptible to human inspection. In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision.
@@
There are various invisible triggers designed for the backdoor attacks on supervised classifiers, such as WaNet [51], ISSBA [41], and filter attack [46]. A straightforward way to achieve imperceptible backdoor attacks in SSL is by directly applying these invisible triggers. However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (Please refer to section 3 and Figure 3 for more details). We then find that this lack of effectiveness is due to the overlapping distributions between the backdoor samples and the augmented samples utilized in SSL. Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (Please refer to section 3 and Figure 3).
@@
Figure 1: Comparison of clean, backdoored samples created by Patch trigger [18], Instagram filter trigger [33], ISSBA trigger [41], WaNet trigger [51] and ours. Their attack success rates tested under the threat model of BadEncoder [30] are beside the name of the methods. Residuals are the difference between clean and backdoored images.
@@
Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in self-supervised learning. In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the self-supervised learning process. we also keep the stealthiness of the trigger by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance. We then implemented our prototype IMPERATIVE (Imperceptible BAckdoor ATtack In Self-superVised LEarning) and tested it on five datasets (i.e., CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50], ImageNet [60]), and across seven classic SSL frameworks (i.e., SimCLR [5], MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3] and CLIP [58], please refer to section 5 and Appendix A) with their various augmentation transforms. The results demonstrate that our method is highly effective and stealthy. In detail, it achieves 99.54% average ASR and 0.95 SSIM. As shown in Figure 1, we compare the backdoor residuals and attack success rates of several methods. Our method exhibits the highest attack success rate while maintaining the highest stealthiness. It also effectively bypasses existing backdoor defenses such as DECREE [15], Beatrix [48], STRIP [16], Grad-CAM [64], Neural Cleanse [83], and commonly studied noises such as JPEG compression, Poisson noise, and Salt&Pepper noise.
@@
Our contributions are summarized as follows: ① We observed that existing imperceptible triggers designed for supervised classifiers have limited effectiveness in SSL. ② We find that the reason behind such ineffectiveness is the coupling feature-space distributions for the backdoor samples and augmented samples in the SSL models. ③ Based on our findings, we propose an imperceptible and effective backdoor attack in SSL by disentangling the distribution of backdoor samples and augmented samples in SSL, while constraining the stealthiness of the triggers during the optimization process. ④ Extensive experiments on five datasets and seven SSL algorithms with their different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
@@ 2.1 Self-Supervised Learning
The objective of SSL is to utilize a vast array of unlabeled data (referred to as the pre-training dataset) to pre-train an image encoder. After the pre-training phase, this encoder can then be used to develop classifiers for various downstream tasks utilizing a smaller set of labeled data (known as the downstream dataset) [71]. SSL typically includes the following pipelines for contrastive learning as outlined in [27, 19]: ① Negative Examples [5, 22, 98, 77, 78]: This approach fosters closeness among positive examples and seeks to maximize the distance between negative examples within the latent space. Notable methods include SimCLRs [5] and MoCo [22]. ② Self-distillation [4, 6, 17, 78]: This strategy involves the use of two identical deep neural networks, termed siamese networks, which share the same architecture but operate with distinct weights. The aim is to enhance the similarity between two differently augmented versions of the same image under certain conditions, such as BYOL [17], SimSiam [6], and DINO [4]. ③ Clustering [3]: This approach applies a clustering mechanism by employing swapped prediction of the representations derived from both encoders in an end-to-end framework, such as SwAV [3]. We will demonstrate that our method is highly effective and stealthy under the SSL algorithms in each category in section 5 and Appendix A.
@@ 2.2 Backdoor Attacks
Numerous strategies for backdoor attacks in SL (supervised-learning) have been introduced [36, 88, 49, 74, 56, 95, 40], which can be classified based on (i) the objectives of the attacks, including attacks tailored to specific inputs [65, 41], specific classes [69], or all inputs [18]; (ii) the methods of attack, such as poisoning the training data [45], identifying susceptible architectures [54], or distributing models with embedded backdoors [28]; and (iii) the criteria for optimization, focusing on the attack’s effectiveness [53], its ability to be transferred [93], or its stealthiness [80, 7, 99, 65]. Since many attacks mentioned above cannot be applied to SSL due to the necessity of labeled data, recent studies have investigated alternative methods for implanting backdoors into models trained with SSL [61, 2, 43, 30]. However, they usually use a patch as a backdoor trigger and are susceptible to human inspection and model simulation.
@@ 2.3 Backdoor Defenses
Various defenses have been developed to defend against backdoor attacks [35, 90, 24, 47, 87, 57, 55, 52, 94, 70]. Most of them are designed to protect against backdoor threats targeting supervised classifiers. These mechanisms work by either preventing attacks during the model training process [85, 25, 23, 79], or by identifying and addressing backdoors in already compromised models in an offline manner [73, 100, 72, 86, 92, 66, 46, 39]. Additionally, some methods are aimed at recognizing backdoor-triggered inputs at the time of inference [8, 11, 16]. DECREE [15] and Beatrix [48] are two strategies specifically designed for SSL, and we will demonstrate that our method can bypass them in subsection 5.4.
3 Observations and Insights
Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in self-supervised learning (SSL). The attack framework for SL and SSL are standard backdoor poisoning [18] and BadEncoder [29], respectively.
@@ 3 Observations and Insights
Figure 3: SimCLR [5] pre-trained model struggled to differentiate between the backdoor samples with WaNet trigger [51] and augmented samples in SimCLR contrastive learning framework.
We first evaluate the effectiveness of existing invisible triggers (i.e., WaNet [51], ISSBA [41], and filter attack [46]) designed for supervised classifiers. In Figure 3, we show their attack success rates on supervised classifiers and self-supervised models. For the experiments on supervised classifiers, the attack method is standard backdoor poisoning [18]. For self-supervised models, we use the attack method introduced in BadEncoder [30] and replace the patch trigger with these invisible triggers. The model used here is ResNet18 [21]. We observed that these invisible triggers, which yield high Attack Success Rates (ASR) in supervised learning, do not perform as effectively in self-supervised learning. In detail, we employed a ResNet18 pre-trained with SimCLR [5] to train and test a binary classification task, aiming to discern whether samples poisoned by WaNet [51] and those augmented through SimCLR’s contrastive learning-based data augmentation are distinguishable. Throughout this process, we maintained consistent feature representations for clean samples between the backdoored and clean models by using utility loss introduced in Jia et al. [32].
Results indicate that the models struggled to differentiate between the two categories. A t-SNE [81] visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement. From this, we infer that the diminished effectiveness of supervised backdoor attack methods in self-supervised learning scenarios may be attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples. Consequently, our method endeavors to disentangle the feature space distribution of these two sets. The experimental results demonstrate that this approach substantially improves ASR in self-supervised learning.
@@ 4.1 Threat Model
Objectives of the Attacker.  The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL (self-supervised learning). We define a backdoored image encoder model as $\mathcal{F}_{\theta}$ and the backdoor injector as $\mathcal{I}_{\phi}$. In this way, a downstream classifier trained based on $\mathcal{F}_{\theta}$, which we define as $\mathcal{C}_{\epsilon}$, could produce a specific prediction $c$ designated by the attacker for inputs $\boldsymbol{x}$ implanted with a trigger chosen by the attacker. The formal definition is shown as follows. $y$ here means the correct label of the input $\boldsymbol{x}$.
@@
The ultimate aim of the attacker is to modify a clean image encoder and create a backdoored version that accomplishes several key objectives: ① Effectiveness: The backdoored model should achieve a high rate of successful attacks while preserving high accuracy in benign conditions. ② Naturalness and Stealthiness: To avoid detection during human inspection, the sample with the trigger should appear authentic and natural. ③ Robustness: The effectiveness of the attack should remain intact even when the triggered samples undergo common pre-processing steps, such as samples under various noises. Furthermore, the attack should bypass prevailing defense mechanisms.
Attacker’s Knowledge and Capabilities.  We assume that the attacker has access to a clean image encoder pre-trained following [29]. Additionally, it is presumed that the attacker has access to a collection of unlabeled images, referred to as shadow dataset. The attacker is also assumed to have access to a few images from the Internet, called reference inputs, for each combination of a target downstream task and a target class. We assume that the attacker is capable of manipulating the training procedure to create an encoder with embedded backdoors. Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in IMPERATIVE to generate stealthy and effective backdoor triggers. However, we assume that the attacker cannot interfere with the training process of these downstream classifiers, such as the training dataset, model framework, and weights.
4.2 Preliminaries
HSV and HSL Color Spaces.  HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing. They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications. They have wide applications in the computer vision domain [62, 76, 42, 13]. More details are shown in subsection D.5.
Wasserstein Distance.  Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71]. Specifically, we focus on the 2-Wasserstein distance, which is elaborated in subsection D.5.

### v2

# Invisible Backdoor Attack against Self-supervised Learning

@@ Abstract
Self-supervised learning (SSL) models are vulnerable to backdoor attacks. Existing backdoor attacks that are effective in SSL often involve noticeable triggers, like colored patches or visible noise, which are vulnerable to human inspection. This paper proposes an imperceptible and effective backdoor attack against self-supervised models. We first find that existing imperceptible triggers designed for supervised learning are less effective in compromising self-supervised models. We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in SSL. Building on this insight, we design an attack using optimized triggers disentangled with the augmented transformation in the SSL, while remaining imperceptible to human vision. Experiments on five datasets and six SSL algorithms demonstrate our attack is highly effective and stealthy. It also has strong resistance to existing backdoor defenses. Our code can be found at https://github.com/Zhang-Henry/INACTIVE.
@@ 1 Introduction
In recent years, Self-Supervised Learning (SSL) has become a powerful approach in deep learning, enabling the learning of rich representations from vast unlabeled data, thus avoiding manual labeling. SSL aims to develop an image encoder that produces similar embeddings for similar images by applying various augmentations to the same image. This pre-trained encoder can be used for different downstream tasks by training compact downstream classifiers with relatively few parameters.
@@
Although SSL has been extensively used in the development of foundational models [6, 20, 5], it is at risk of backdoor attacks [27, 56, 34, 68], where the attacker embeds hidden malicious behavior within the encoder. The backdoor can be inherited to the downstream task. The downstream classifier predicts a specific target label if the input contains a pre-defined backdoor trigger. Existing backdoor attacks on SSL such as BadEncoder [27] achieve high attack success rates (ASR). However, a common drawback of these effective attacks is that their trigger patterns are obvious, making them susceptible to human inspection. Moreover, while data-poisoning-based attacks CTRL [34] and BLTO [63] are relatively stealthy, their ASRs are suboptimal, For example, on CIFAR10 CTRL only has 61.90% ASR under BYOL framework and BLTO only has 84.63% ASR under SimSiam framework. Furthermore, they also rely on the downstream dataset matching the pre-training dataset distribution, limiting effectiveness across diverse datasets. In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision without this distribution dependency.
@@
There are various invisible triggers designed for the backdoor attacks on supervised classifiers, such as WaNet [47], ISSBA [36], and filter attack [42]. A straightforward way to achieve imperceptible backdoor attacks in SSL is by directly applying these invisible triggers. However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (see Figure 2). We then find that this lack of effectiveness is due to the overlapping distributions between the backdoor samples and the augmented samples utilized in SSL. Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (see Figure 3).
@@
Figure 1: Comparison of clean, backdoored samples created by Patch trigger used by BadEncoder [27] and DRUPE [68], Instagram filter trigger [30], ISSBA trigger [36], WaNet trigger [47] and ours. Except for DRUPE, the ASRs are tested under the threat model of BadEncoder. Residuals are the difference between clean and backdoored images. Our method achieves the highest ASR while maintaining trigger stealthiness, while other methods either have a much lower ASR or use more easily detectable triggers.
Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in SSL. The attack framework for SL and SSL are standard backdoor poisoning [17] and BadEncoder [27], respectively.
@@
Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in SSL. In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the SSL process. We also keep the trigger stealthy by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance. We then implemented our prototype INACTIVE (INvisible bACkdoor aTtack In self-superVised lEarning) and tested it on five datasets (i.e., CIFAR10, STL10, GTSRB, SVHN, ImageNet), and across six classic SSL frameworks (i.e., SimCLR [6], MoCo [20], BYOL [16], SimSiam [7], SwAV [4] and CLIP [53] (See Table 1 and Table A1) with their various augmentation transforms (See Table A5). The results demonstrate that our method is highly effective and stealthy. In detail, it achieves an average of 99.09% ASR (See Table 1), 0.9763 SSIM, 41.07 PSNR, 0.0046 LIPIS, 0.9751 FSIM, and 13.281 FID (See Table A9). As shown in Figure 1, we compare several methods’ backdoor residuals and ASRs. Our method exhibits the highest ASR while maintaining the highest stealthiness. It also effectively bypasses existing backdoor defenses such as DECREE [13], Beatrix [44], ASSET [48], STRIP [14], Grad-CAM [58], Neural Cleanse [72], and various noise, i.e., JPEG compression, Poisson noise, and Salt&Pepper noise.
@@
Our contributions are summarized as follows: ① We observed that existing imperceptible triggers designed for supervised classifiers have limited effectiveness in SSL. ② We find that the reason behind such ineffectiveness is the coupling feature-space distributions for the backdoor samples and augmented samples in the SSL models. ③ Based on our findings, we propose an imperceptible and effective backdoor attack in SSL by disentangling the distribution of backdoor samples and augmented samples in SSL, while constraining the stealthiness of the triggers during the optimization process. ④ Extensive experiments on five datasets and six SSL algorithms with different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
@@ 2.1 Self-Supervised Learning
The goal of SSL is to leverage a large amount of unlabeled data in the pre-training dataset to pre-train an image encoder, which can then be used to create classifiers for various downstream tasks with a smaller set of labeled data [68]. SSL pipelines for contrastive learning typically include the following approaches [26, 18]: ① Negative Examples: Promotes proximity among positive examples while maximizing the distance between negative examples in the latent space, as seen in SimCLRs [6] and MoCo [20]. ② Self-distillation: Utilizes two identical Siamese networks with different weights to increase the similarity between differently augmented versions of the same image, such as in BYOL [16] and SimSiam [7]. ③ Clustering: Implements a clustering mechanism with swapped prediction of representations from both encoders, as in SwAV [4]. Our method is shown to be highly effective and stealthy under the SSL algorithms in Table 1 and Table A1.
@@ 2.2 Backdoor Attacks
Backdoor attacks were initially proposed for supervised learning (SL) to modify a model’s behavior on specific inputs or classes while keeping its general performance intact [33, 77, 45, 67, 85, 51, 84, 15, 2]. Early backdoor attacks commonly utilized visible triggers like distinctive patches that are easily detectable through visual inspection [17, 59, 64]. To enhance stealth, subsequent research introduced invisible triggers, which are subtle and blend into the background, helping these attacks evade both human inspection and certain automated defenses [36, 47].
Since many of these attacks rely on labeled data, recent studies have explored alternative backdoor implantation techniques in SSL models [56, 3, 39, 27, 63]. However, these typically use a visible backdoor trigger, such as a patch, making them prone to human detection and model simulation. The advantage of invisible triggers in SSL is clear: they improve attack stealth, bypassing some conventional defenses that focus on detecting visible anomalies [13, 69]. However, as we will demonstrate, directly applying existing invisible triggers designed for SL to SSL tasks results in limited attack effectiveness. Moreover, backdoor attacks in SSL are generally divided into two types: training-time backdoor injection attacks like BadEncoder [27], which require control of the training in the backdoor injection process, and data-poisoning-based attacks like CTRL [34] and BLTO [63], which rely on poisoned data without needing model specifics. BadEncoder modifies a pre-trained encoder to embed triggers that cause targeted misclassifications in downstream tasks. By aligning the features of triggered images with an attacker-chosen class, downstream classifiers misclassify triggered inputs while maintaining accuracy on clean data. Although CTRL and BLTO are relatively stealthy, they achieve lower ASRs than our method and depend on the downstream dataset matching the pre-training distribution, limiting their versatility across diverse datasets. In this paper, we focus on training-time backdoor injection attack due to it has higher attack effectiveness and transferability.
@@ 2.3 Backdoor Defenses
Various defenses have been developed against backdoor attacks [32, 78, 23, 43, 76, 52, 50, 49, 83, 8, 81], primarily targeting supervised classifiers. These defenses either prevent attacks during training [74, 24, 22, 70] or detect and mitigate backdoors in compromised models offline [66, 89, 65, 75, 80, 60, 42, 37]. Some methods also detect backdoor-triggered inputs during inference [14]. Defense methods like DECREE [13], Beatrix [44], and ASSET [48] are designed for SSL, primarily relying on the visible characteristics of triggers to detect backdoors. In contrast, our method uses invisible triggers, effectively bypassing these defenses by breaking their reliance on visual anomalies and making detection more challenging.
3 Observations and Analysis
@@
Figure 3: t-SNE visualization of the feature space in the inherent augmentation and backdoor trigger space. The SimCLR [6] pre-trained model struggled to differentiate between backdoor samples injected with the WaNet trigger [47] and the augmented samples within the SimCLR contrastive learning framework.
Invisible Trigger Designed for SL Fails in SSL. We first assess the effectiveness of existing invisible triggers (WaNet, ISSBA, and filter attack) designed for supervised classifiers. Figure 2 displays their ASRs on both supervised classifiers and self-supervised models. For supervised classifiers, the standard backdoor poisoning method [17] is used. For self-supervised models, we apply the BadEncoder method, replacing the patch trigger with these invisible triggers, using ResNet18. We find that these triggers, which achieve high ASR in supervised learning, are less effective in SSL. We then investigate the underlying reason for these results and focus on the following research question: Why does the effective invisible backdoor trigger designed for supervised learning fail on self-supervised learning?
Cause of the Failure: Entanglement of the Inherent Augmentations and Backdoor Trigger. We find that the entanglement of the inherent augmentations in contrastive learning can cause the failure of the backdoor injection with such triggers. We provide our analysis in this section. One of the core training losses of contrastive learning can be formulated as maximizing the feature space similarity between the augmented samples modified from the same training samples:
$\arg\max_{\theta_{\mathcal{F}}}s(\mathcal{F}_{\theta}(\mathcal{A}_{1}(\bm{x})),\mathcal{F}_{\theta}(\mathcal{A}_{2}(\bm{x})))$ (1)
where $s(\cdot,\cdot)$ denotes the similarity measurement, $\mathcal{F}_{\theta}$ is the encoder in training ($\theta_{\mathcal{F}}$ is its parameters), $\bm{x}$ is the training sample, $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are different augmentations sampled from the predefined augmentation space $\mathcal{S}_{\mathcal{A}}$. Different from the predefined augmentation space, we also define the learned augmentation space for trained encoders $\mathcal{S}_{\mathcal{A}}^{\prime}$ as the space including a set of transformations where any pair within it can achieve high pairwise similarity on augmented versions of the same sample when processed by the trained encoder. We also define perfectly-trained encoder is the encoder that achieves maximal similarity described in Equation 1 for all samples and all possible transformations used, i.e., $s(\mathcal{F}_{\theta}(\mathcal{A}_{1}(\bm{x})),\mathcal{F}_{\theta}(\mathcal{A}_{2}(\bm{x})))=1,\forall\bm{x}\in\mathcal{X},\forall\mathcal{A}_{1},\mathcal{A}_{2}\in\mathcal{S}_{\mathcal{A}}$, where $\mathcal{X}$ is the input space. Based on this, we have the following theorem:
Theorem 3.1.
Given a perfectly-trained encoder $\mathcal{F}_{\theta}$ based on the augmentations sampled from predefined augmentation space $\mathcal{S}_{\mathcal{A}}$, it is impossible to inject a backdoor with trigger function $\mathcal{I}\in\mathcal{S}_{\mathcal{A}}$.
The proof of this theorem can be found in the section A4. In practice, the boundary of the learned augmentation space for trained encoders $\mathcal{S}_{\mathcal{A}}^{\prime}$ is often imprecise, and it potentially reflects a relaxation of the predefined augmentation space. Consequently, using trigger functions that are not precisely within the predefined augmentation space $\mathcal{S}_{\mathcal{A}}$ but are instead distributionally close to it can also make achieving high attack success rates hard.
Empirical Evidence. We also conduct experiments to confirm the invisible triggers designed for supervised learning are actually entangled with the inherent augmentations in self-supervised learning. Specifically, we use a ResNet18 pre-trained with SimCLR for a binary classification task to differentiate between samples poisoned by WaNet and those augmented by SimCLR. We ensure consistent feature representations for clean samples between the backdoored and clean models using utility loss from Jia et al [29]. Results indicate that the models struggle to differentiate between the two categories. A t-SNE visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement. From this, we infer that the diminished effectiveness of supervised backdoor attack methods in SSL scenarios is attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples. The reason for the entanglement phenomenon on the invisible backdoor trigger is that the learned augmentation space reflects a relaxation of the predefined augmentation space, and such relaxation covers most of the invisible minor transformations. Thus, we aim to search for invisible transformations that can escape the inherent augmentation space.
@@ 4.1 Threat Model
We follow the well-defined training-based backdoor injection threat model introduced in BadEncoder [27].
Objectives of the Attacker. The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL. We define a backdoored image encoder model as $\mathcal{F}_{\theta}$ and the backdoor injector as $\mathcal{I}_{\phi}$. In this way, a downstream classifier trained based on $\mathcal{F}_{\theta}$, which we define as $\mathcal{C}_{\epsilon}$, could produce a specific prediction $c$ designated by the attacker for inputs $\boldsymbol{x}$ implanted with a trigger chosen by the attacker. The formal definition is shown as follows. $y$ here means the correct label of the input $\boldsymbol{x}$.
@@
The attacker’s goal is to modify a clean image encoder to create a backdoored version that meets two key objectives: ① Effectiveness: The backdoored model should maintain a high attack success rate while preserving accuracy in benign conditions, keeping backdoored accuracy close to clean accuracy for downstream classifiers. ② Naturalness and Stealthiness: The triggered sample should appear authentic and natural to avoid detection by human inspection.
Attacker’s Knowledge and Capabilities. Following BadEncoder [27], we assume that the attacker has access to a pre-trained clean image encoder and the attacker has full knowledge about the pre-trained encoder, such as the SSL method and the detailed contrastive augmentation operation used in pre-training. Additionally, it is presumed that the attacker can access a collection of unlabeled images, referred to as shadow dataset. The attacker is also assumed to have access to a few images from the Internet, called reference inputs, for each combination of a target downstream task and a target class. We assume that the attacker can manipulate the training procedure to create an encoder with embedded backdoors. Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in INACTIVE to generate stealthy and effective backdoor triggers. However, we assume that the attacker cannot interfere with the training process of these downstream classifiers, such as the training dataset, model framework, and weights. Unlike data-poisoning-based methods such as CTRL [34] and BLTO [63], our approach does not rely on matching distributions between pre-training and downstream datasets, allowing for broader applicability without interfering in downstream training data, models, or weights.